OpenStudy (anonymous):

Lyle has a bowl of fruit on his dining room table. It contains 2 oranges, 6 apples, and 2 peaches. Suppose he selects a piece of fruit at random from the bowl. What is the probability that he selects either an orange or a peach?

OpenStudy (amistre64):

we add the probs if i recall correctly

OpenStudy (amistre64):

4 options out of 10

OpenStudy (amistre64):

4/10 = 2/5, and i cant recall if that needs to be decimated or not

OpenStudy (anonymous):

what amistre said \[\frac{4}{10}=\frac{2}{5}=.4\] which ever form you like best

OpenStudy (anonymous):

A.20% B.40% C.60% D.80%

OpenStudy (anonymous):

oooooooooooh another form too. wow

OpenStudy (amistre64):

.4 is 40% :)

OpenStudy (amistre64):

probability is not a percentage tho is it?

OpenStudy (amistre64):

my stats teacher would always get us with that ...

OpenStudy (anonymous):

OK